# Visual Manipulation with Legs

### v1

@@ 2 Related Work
Quadrupedal Locomotion: Legged robots have traditionally relied on proprioceptive senses for navigating complex terrains, using model-based methods that ensure stability but require significant computational power [11, 12, 13, 14, 15, 16, 17, 18]. Recent advancements in model-free reinforcement learning have reduced computational demands while enhancing generalization [19, 20, 21, 22, 23, 24, 25]. Although these techniques have primarily been applied to locomotion, recent work has integrated exteroceptive perception, such as elevation maps and depth sensing, to enhance high-level planning [26, 27, 28, 29, 30, 31, 32, 33, 34]. While most research uses these approaches for locomotion, our work uniquely applies them to dexterous manipulation, leveraging model-based controllers for precise end-effector positional control.
@@
Non-Prehensile Manipulation: Non-prehensile manipulation, which involves interacting with objects without grasping, has been a focus for tasks like pushing, sliding, and flipping. Foundational work by Mason [4] established the mechanics of these tasks, while more recent approaches have integrated reinforcement learning and model-based methods to enhance precision and adaptability [37, 38]. For instance, Zhou et al. [39] developed learning frameworks for robust manipulation in unstructured environments. Building on these advances, our work applies non-prehensile techniques in legged robots, achieving 6D object control through precise actions like pushing and flipping.
Manipulation with Legged Robot: Legged robots are increasingly recognized for their potential in manipulation tasks. Leveraging research that integrates leg and arm mechanisms [40] and biomimetic designs, these robots enhance movement and object transportation [41]. Existing methods predominantly use mounted arms [1, 2, 42] or the robot’s body [5, 6] for basic tasks, with few employing legs. In this context, some works have extended quadruped capabilities by attaching small grippers at the end-effectors of their legs to facilitate prehensile skills  [43]. Others have utilized the legs for broader, large-scale manipulations, such as door pushing and basket lifting [3, 44, 45, 9] or rotating large yoga balls [46] but sacrificing mobility. In contrast, our approach emphasizes non-prehensile manipulation, using the quadruped’s legs to directly control objects with precise actions like pushing and flipping smaller items, achieving 6D object control without relying on additional appendages and maintaining full mobility.

### v3

@@ 2 Related Work
Legged Robot Locomotion: Legged robots have traditionally relied on proprioceptive senses for navigating complex terrains, using model-based methods that ensure stability but require significant computational power [11, 12, 13, 14, 15, 16, 17, 18]. Recent advancements in model-free reinforcement learning have reduced computational demands while enhancing generalization [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Although these techniques have primarily been applied to locomotion, recent work has integrated exteroceptive perception, such as elevation maps and depth sensing, to enhance high-level planning [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. While most research uses these approaches for locomotion, our work uniquely applies them to dexterous manipulation, leveraging model-based controllers for precise end-effector positional control.
@@
Non-Prehensile Manipulation: Non-prehensile manipulation, which involves interacting with objects without grasping, has been a focus for tasks like pushing, sliding, and flipping. Foundational work by Mason [4] established the mechanics of these tasks, while more recent approaches have integrated reinforcement learning and model-based methods to enhance precision and adaptability [43, 44, 45]. For instance, Zhou et al. [46] developed learning frameworks for robust manipulation in unstructured environments. Building on these advances, our work applies non-prehensile techniques in legged robots, achieving 6D object control through precise actions like pushing and flipping.
Manipulation with Legged Robot: Legged robots are increasingly recognized for their potential in manipulation tasks. Leveraging research that integrates leg and arm mechanisms [47] and biomimetic designs, these robots enhance movement and object transportation [48]. Existing methods predominantly use mounted arms [1, 2, 49, 50, 51] or the robot’s body [5, 6, 52, 53, 54] for basic tasks, with few employing legs. In this context, some works have extended quadruped capabilities by attaching small grippers at the end-effectors of their legs to facilitate prehensile skills  [55]. Others have utilized the legs for broader, large-scale manipulations, such as door pushing and basket lifting [3, 56, 57, 9] or rotating large yoga balls [58] but sacrificing mobility. In contrast, our approach emphasizes non-prehensile manipulation, using the quadruped’s legs to directly control objects with precise actions like pushing and flipping smaller items, achieving 6D object control without relying on additional appendages and maintaining full mobility.